OpenStudy (anonymous):

the limit as n -> infinity sqrt(n)[(sqrt(9n+1)-sqrt(9n)].

OpenStudy (anonymous):

looks as though that diverges unless there is supposed to be a division in that formula

OpenStudy (anonymous):

using the product rule for limits you can first evaluate the limit of sqrt(n) and then the limit in the bigger brackets

OpenStudy (anonymous):

1/6 seems a very peculiar answer

OpenStudy (anonymous):

there is nothing to stop that limit diverging

OpenStudy (nikvist):

\[\lim_{n\rightarrow +\infty}\sqrt{n}\cdot (\sqrt{9n+1}-\sqrt{9n})=\] \[=\lim_{n\rightarrow +\infty}\sqrt{n}\cdot (\sqrt{9n+1}-\sqrt{9n}) \cdot\frac{\sqrt{9n+1}+\sqrt{9n}}{\sqrt{9n+1}+\sqrt{9n}}=\] \[=\lim_{n\rightarrow +\infty}\frac{\sqrt{n}}{\sqrt{9n+1}+\sqrt{9n}}=\lim_{n\rightarrow +\infty}\frac{1}{\sqrt{9+1/n}+\sqrt{9}}=\frac{1}{6}\]

OpenStudy (anonymous):

thanks